# Tiling spaces are covering spaces over irrational tori

### v1

@@ 4 Diffeological and sturmian equivalences for Tiling spaces
Then each orbit in $X_{\alpha}$ is projected onto a class of $\mathbb{T}_{\alpha}$. But recall that there is a double orbit corresponding to sequences differing in two consecutive letters. Including both of these orbits, we obtain a $\mathbb{Z}$-principal fiber bundle $X_{\alpha}\to\mathbb{T}_{\alpha}$ with a singular point with a double fiber. In order to avoid this singularity, we consider the double pointed irrational torus
$\mathbb{T}^{\bullet}_{\alpha}=\mathbb{T}_{\alpha}\times\{0,1\}/\sim,\quad\text{with}\quad([x],j)\sim([y],i)\quad\text{iff}\quad[x]=[y]\quad\text{and}\quad[x]\neq[0].$ (19)
Where $[x]$ is the equivalent class of $x\in\mathbb{R}$ through the projection $\mathrm{pr}_{\alpha}$ (7). Now we can extend through $\Psi$ to obtain the conmutative diagram of smooth maps (it should be clear by now that the contruction of $\mathbb{T}_{\alpha}^{\bullet}$ is smooth)
${X_{\alpha}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\Psi}$$\scriptstyle{p}$${\Omega_{\alpha}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\pi_{\alpha}^{\bullet}}$${\mathbb{T}_{\alpha}^{\bullet}.}$ (20)
We have seen already in Lemma 2 and Section 2.2.3 that the Kronecker flow $\mathcal{K}_{\alpha}$ is a $\mathbb{R}$-principal bundle over $\mathbb{T}_{\alpha}$. So there is only left to prove that the new action map $\mathcal{A}^{\mathbb{R}}_{\Omega_{\alpha}}$ is still an induction. But by construction the extra fiber of $\Omega_{\alpha}$ is projected onto the extra point of $\mathbb{T}_{\alpha}^{\bullet}$, so if $\mathcal{A}^{\mathbb{R}}_{\mathcal{K}_{\alpha}}$ is an induction then $\mathcal{A}^{\mathbb{R}}_{\Omega_{\alpha}}$ must also be. Thus the surjection $\pi_{\alpha}^{\bullet}$ is $\mathbb{R}$-principal fibration.
An importat consequence of these results is that we get to define a fundamental group for $\Omega$.
@@
$\pi_{1}(\Omega)=\mathbb{Z}+\mathbb{Z}+\alpha\mathbb{Z}.$
@@
The proof relies on relative homotopy and the long exact sequence of homotopy of a pair [21, Postulate 5.19] – we omit here the basepoint because we are dealing with homogeneus spaces. We consider the relative homotopy of the pair $\mathcal{K}_{\alpha}\subset\Omega_{\alpha}$:
$\ldots\longrightarrow\pi_{2}(\Omega_{\alpha},K_{\alpha})\longrightarrow\pi_{1}(\mathcal{K}_{\alpha})\longrightarrow\pi_{1}(\Omega_{\alpha})\longrightarrow\pi_{1}(\Omega_{\alpha},\mathcal{K}_{\alpha})\longrightarrow\pi_{0}(\mathcal{K}_{\alpha})\longrightarrow\pi_{0}(\Omega_{\alpha})$
We can straightforward check that $\pi_{2}(\Omega_{\alpha},\mathcal{K}_{\alpha})=0$, because there is no room for 2-dimensional paths; also $\pi_{1}(\mathcal{K}_{\alpha})=\mathbb{Z}+\alpha\mathbb{Z}$, see the exact sequence 13. As $\pi_{1}(\Omega_{\alpha},\mathcal{K}_{\alpha})$ is isomorphic to the fundamental group of the line with double origin, which is $\mathbb{Z}$. And $\pi_{0}(\mathcal{K}_{\alpha})=0$ because it is connected.
This means that
$\pi_{1}(\Omega_{\alpha})=\pi_{1}(\mathcal{K}_{\alpha})+\pi_{1}(\Omega_{\alpha},\mathcal{K}_{\alpha})=\mathbb{Z}+\mathbb{Z}+\alpha\mathbb{Z}$
@@
If $X_{\alpha}$ and $X_{\beta}$ are strong orbit equivalent there is a cocycle map $n:X_{\alpha}\to\mathbb{Z}$ and a diffeomorphism $h:X_{\alpha}\to X_{\beta}$ such that $h(\sigma(S))=\sigma^{n(S)}(h(S))$. Then, it is possible lift $h$ through the smooth maps $\hat{I}_{\bullet},\Psi_{\alpha}$ and $\Psi_{\beta}$ so as to obtain the following commutative diagram.
@@
The lift $\tilde{h}:\Omega_{\alpha}\to\Omega_{\beta}$ is more than an orbital equivalence. To see this, we define another cocycle $m:X_{\alpha}\times\mathbb{R}\to\mathbb{Z}$ such that $\sigma^{m(T,x)}(\Phi_{\alpha}(T))=\Phi_{\alpha}(T+x)$. Then we have the following relation
@@
with $S=\Phi_{\alpha}(T)$ and $d$ the distance function of the metric of $\Omega_{\beta}$, and $\delta_{\beta}$ as in 17.
@@
The equivalence $i)\Leftrightarrow ii)$ is by definition, $ii)\Leftrightarrow iv)$ is Lemma 1 and $iii)\Leftrightarrow iv)$ Lemma 3.
Our results provide a classification of one-dimensional tiling spaces as conjugacy classes of irrational numbers under the modular group $\mathrm{GL}(2,\mathbb{Z})$, which implies what we have called strong orbit equivalence for tiling spaces. Our last result shows that there is a special representative within this classes which is a substitution tiling space.
Theorem D.
@@
Recall from Remark 3 and results from [27, 16] that a cut-and-project tiling admit a substitution if the irrational slope $\alpha$ is a quadratic Pisot number. Moreover, a theorem by Galois [14] asserts that quadratic Pisot numbers have purely periodic continued fraction. This, together with the well known fact that the continued fraction of conjugates under the modular group differ in a finite number of coeficients, implies that there is a Pisot number within this class of conjugation, provided the irrational number $\alpha$ is quadratic. This Pisot number is expresed as the continued fraction of the periodic tail of $\alpha$.
@@ Acknowledgments
We thank Ana Rechtman, Alberto Verjovsky, Ángel Zaldívar, José Aliste-Prieto, Andrés Navas and Fabien Durand for their support in different stages of this research.

### v3

@@ 4 Diffeological and sturmian equivalences for Tiling spaces
Then each orbit in $X_{\alpha}$ is projected onto a class of $\mathbb{T}_{\alpha}$. The mapping is smooth because so is $\mathrm{pr}_{\alpha}$.
But recall from Remark 1 that $\Omega_{\alpha}$ contains an additional orbit corresponding to sequences differing in two consecutive letters. By considering both of these orbits, we obtain a $\mathbb{Z}$-principal fiber bundle $X_{\alpha}\to\mathbb{T}_{\alpha}$ with a singular point with a double fiber. We deal with this singularity in the following way.
Note that the equivalence classes $[x]$ of $\mathbb{T}^{2}/\mathcal{S}_{\alpha}$ correspond to orbits under the action by translation on $\Omega_{\alpha}$, so they are path connected and diffeomorphic to $\mathbb{R}$. We define the irrational tori with two origins $\mathbb{T}_{\alpha}^{\bullet}$ as the quotient of the punctured torus $\mathbb{T}^{2}-\{e\}$ by the equivalence relation $x\sim y$ if and only if $[x]=[y]$ and both $x$ and $y$ are in the same path connected component of $[x]-\{e\}$, for $e\in\mathcal{S}_{\alpha}$ the identity element of the group $\mathbb{T}^{2}$. The class $\mathcal{S}_{\alpha}$ is divided in two rays, which are diffeomorphic to the real line.
Once this construction is made, it is straightforward to check that the quotient map $\pi:\mathbb{T}^{2}-\{e\}\to\mathbb{T}_{\alpha}^{\bullet}$ has fibrating structure groupoid $K_{\pi}$. So it is a diffeological fibration with fiber $\mathbb{R}$. The inclusion $i:\mathbb{T}^{2}-\{e\}\to\mathbb{T}^{2}$ induces the following conmutative diagram:
[…]
note that $I_{\bullet}$ is surjective. Thus, we define the $\mathbb{Z}$-principal fiber bundle $X_{\alpha}\to\mathbb{T}_{\alpha}^{\bullet}$
Now we can extend through $\Psi$ in order to obtain the following conmutative diagram of smooth maps
[…]
We have seen already in Lemma 2 and Section 2.2.3 that the Kronecker flow $\mathcal{K}_{\alpha}$ is a $\mathbb{R}$-principal bundle over $\mathbb{T}_{\alpha}$ and that there is a bijection between $\Omega_{\alpha}-\{O(T_{1})\}$ and $\mathcal{K}_{\alpha}$. So there is only left to prove that the new action map $\mathcal{A}^{\mathbb{R}}_{\Omega_{\alpha}}$ is still an induction. But by construction, the additional fiber of $\Omega_{\alpha}$ is smoothly projected onto the extra point of $\mathbb{T}_{\alpha}^{\bullet}$, so if $\mathcal{A}^{\mathbb{R}}_{\mathcal{K}_{\alpha}}$ is an induction then $\mathcal{A}^{\mathbb{R}}_{\Omega_{\alpha}}$ must also be. Thus the surjection $\pi_{\alpha}^{\bullet}$ is $\mathbb{R}$-principal fibration.
As a consequence, we are able to compute the fundamental group of $\Omega_{\alpha}$.
@@
$\pi_{1}(\Omega_{\alpha})=\mathbb{F}_{2},$
where $\mathbb{F}_{2}$ is the free group generated by two elements.
@@
The proof relies on the exact sequence of homotopy [19, Postulate 8.21] of the diffeological fibration $\pi:\mathbb{T}^{2}-\{e\}\to\mathbb{T}_{\alpha}^{\bullet}$,
$\ldots\longrightarrow\pi_{2}(\mathbb{T}_{\alpha}^{\bullet})\longrightarrow\pi_{1}(\mathbb{R})\longrightarrow\pi_{1}(\mathbb{T}^{2}-\{e\})\longrightarrow\pi_{1}(\mathbb{T}^{\bullet}_{\alpha})\longrightarrow\pi_{0}(\mathbb{R})\longrightarrow\pi_{0}(\mathbb{T}^{2}-\{e\})$
As $\mathbb{R}$ is simply connected, we have that the fundamental groups $\pi_{1}(\mathbb{T}^{\bullet}_{\alpha}),\pi_{1}(\mathbb{T}^{2}-\{e\})$ are isomorphic. Note that $\Omega_{\alpha}$ is $\mathbb{R}$-principal bundle on $\mathbb{T}^{\bullet}_{\alpha}$, thus applying the exact sequence of homotopy we obtain that
$\pi_{1}(\Omega_{\alpha})=\pi_{1}(\mathbb{T}^{\bullet}_{\alpha})=\pi_{1}(\mathbb{T}^{2}-\{e\})=\mathbb{F}_{2}.$
@@
If $X_{\alpha}$ and $X_{\beta}$ are strong orbit equivalent, there is a cocycle map $n:X_{\alpha}\to\mathbb{Z}$ and a diffeomorphism $h:X_{\alpha}\to X_{\beta}$ such that $h(\sigma(S))=\sigma^{n(S)}(h(S))$. Then it is possible lift $h$ through the smooth maps $\hat{I}_{\bullet},\Psi_{\alpha}$ and $\Psi_{\beta}$ so as to obtain the following commutative diagram.
@@
The lift $\tilde{h}:\Omega_{\alpha}\to\Omega_{\beta}$ is more than an orbital equivalence. To see this, we define another cocycle $m:\Omega_{\alpha}\times\mathbb{R}\to\mathbb{Z}$ such that $\sigma^{m(T,x)}(\Phi_{\alpha}(T))=\Phi_{\alpha}(T+x)$. Then we have the following relation
@@
where $S=\Phi_{\alpha}(T)$, $d$ is the distance function of the metric of $\Omega_{\beta}$, and $\delta_{\beta}$ as defined in 17.
@@
The equivalence $i)\Leftrightarrow ii)$ is by definition, $ii)\Leftrightarrow iv)$ is Lemma 1, and $iii)\Leftrightarrow iv)$ is Lemma 3.
Our results provide a classification of one-dimensional tiling spaces as conjugacy classes of irrational numbers under the modular group $\mathrm{GL}(2,\mathbb{Z})$, which implies what we have called strong orbit equivalence for tiling spaces. This turns out to be an alternative approach to some old results about substitution sturmian sequences [12, Chapter 6]. In the present context they are stated as
Corollary 2.
@@
It is known from [12, 29, 17] that a cut-and-project tiling admit a substitution if the irrational slope $\alpha$ is a quadratic Pisot number. Moreover, a theorem by Galois [15] asserts that quadratic Pisot numbers have purely periodic continued fraction. This, together with the well known fact that the continued fraction of conjugates under the modular group differ in a finite number of coeficients, implies that there is a Pisot number within this class of conjugation, provided the irrational number $\alpha$ is quadratic. This Pisot number is expresed as the continued fraction of the periodic tail of $\alpha$.
@@ Acknowledgments
We thank Ana Rechtman, Alberto Verjovsky, Ángel Zaldívar, José Aliste-Prieto, Andrés Navas, Fabien Durand and Bernardo Villarreal for their support and advice during different stages of this research.